Evaluate the ∫∫s y dS where S is the portion of the cylinder x2 + y2 = 3 that lies between z = 0 and z = 3.
1
Expert's answer
2020-04-28T16:52:36-0400

Consider the surface "S" , which is the portion of the cylinder "x^2+y^2=3" lies between the planes "z=0" and "z=3" .


Let "x=\sqrt{3}\cos u" , "y=\sqrt{3}\sin u" and "z=v" , where "0\leq u \leq2\pi, 0 \leq v \leq 3"


So, the parametric equation of surface can be represented as,



"r(u,v)=⟨\sqrt{3}\cos u,\sqrt{3}\sin u,v ⟩"


The normal to thee surface "S" is,



"r_u×r_v=\begin{Vmatrix}
 i & j & k \\
 -\sqrt{3}\sin(u) & \sqrt{3}\cos(u) & 0\\
0 & 0 & 1
\end{Vmatrix}
=⟨\sqrt{3}\cos(u),\sqrt{3}\sin(u),0 ⟩"

Now, find the surface integral as,


"\iint_S(y)dS=\iint_{uv}(\sqrt{3}\sin u)∥ r_u×r_v∥dudv"


"=\int_{0}^{2\pi}\int_{0}^{3}\sqrt{3}\sin u \sqrt{(\sqrt{3}\sin u)^2+(\sqrt{3}\cos u)^2}dvdu"

"=\int_{0}^{2\pi}\int_{0}^{3}\sqrt{3}\sin u (\sqrt{3})dvdu"

"=\int_{0}^{2\pi}\int_{0}^{3}3\sin u dvdu"

"=3[-\cos u ]_{0}^{2\pi}[ v]_{0}^{3}"

"=-3(1-1)(3-0)"

"=-3(0)(3)"

"=0"


Therefore, the surface integral is "\iint_S(y)dS=0"